OpenStudy (katielong):

find the equation of the line that is perpendicular to the line y=4x-2 and passes through the point (8,1). thanks =)

OpenStudy (kymber):

Perpendicular lines have opposite reciprocal slopes plug in what you know into y = mx + b & solve for b

OpenStudy (katielong):

so its y=-1/4x + ?

OpenStudy (kymber):

Yup, plug ( 8, 1) into ( x, y ) and solve for b

OpenStudy (katielong):

1=-2+? ?=3... so y=-1/4+3...?

OpenStudy (kymber):

Yeah! :D

OpenStudy (katielong):

woooo! cheers dudee :D